OpenStudy (anonymous):

Find the angle of intersection of the following pairs of curves: y=2x, x^5+ y^5=33

myininaya (myininaya):

when i did the following: x^5+(2x)^5=33 33x^5=33 x^5=1 x=1 => y=2(1)=2 intersection occurs at (1,2) did you get this far?

OpenStudy (anonymous):

yes..

OpenStudy (anonymous):

myininaya pls help me after u done this http://openstudy.com/groups/mathematics/updates/4e94435d0b8b7d552ae41420 sry for interupting

myininaya (myininaya):

i wonder which angle you want because you have 4 angles

myininaya (myininaya):

the 4 angles that was created by the intersection

myininaya (myininaya):

i know how to find the angle between two vectors ....

OpenStudy (anonymous):

Actually the answer was 67.0 degress.. I just can't get the right solution..

myininaya (myininaya):

maybe i should try converting to a polar equation

myininaya (myininaya):

hmmm... let me see what happens

myininaya (myininaya):

oh wait that might work...

myininaya (myininaya):

so we have x=1 right? this implies rcos(theta)=1

OpenStudy (anonymous):

yup yup

myininaya (myininaya):

but r=sqrt{x^2+y^2}=sqrt{1^2+2^2}=sqrt{1+4}=sqrt{5}

myininaya (myininaya):

so we have \[\cos(\theta)=\frac{1}{\sqrt{5}} => \theta=\cos^{-1}(\frac{1}{\sqrt{5}})\]

myininaya (myininaya):

approzimately is 63.535 degrees

myininaya (myininaya):

oops 63.435

myininaya (myininaya):

i like my answer it makes sense

OpenStudy (anonymous):

Thank you.

myininaya (myininaya):

its pretty close to 67 too

myininaya (myininaya):

not too far off

OpenStudy (anonymous):

yep.. i think i got it.. :)

myininaya (myininaya):

cool

myininaya (myininaya):

good question